# InverseMatrixVT3D: An Efficient Projection Matrix-Based Approach for 3D Occupancy Prediction

### v1

@@ IV-A Implementation Details
The InverseMatrixVT3D uses ResNet101-DCN [54, 55] as a 2D backbone with a checkpoint from FCOS3D [56] to extract image features. The features of stages 1,2,3 of the backbone are fed to FPN [57], resulting in 3-level multi-scale image features. The network architecture comprises four levels ($L=4$), with no skip connection applied to the highest level. For the paths corresponding to levels 1, 2, and 3, we employ divided schemas of N=3, 4, and 5, respectively, to create sets of sample points. The AdamW optimizer with an initial learning rate of 5e-5 and weight decay of 0.01 is employed for optimization. The learning rate is decayed using a multi-step scheduler. For the 3D semantic occupancy prediction task, following the approach presented in [39, 37], we incorporate focal loss [58], Lovasz-softmax loss [59] and scene-class affinity loss [37] to handle the significant sparsity of free space in the ground truth dataset. Regarding data augmentation, random resize, rotation and flip operations are applied in the image space, following established practices for BEV-based 3D object detection [23, 24, 25, 26] and the compared methods [38, 39, 40, 37]. Our model includes two resolution settings. The predicted occupancy for evaluation is specified at a resolution of $256\times 256\times 32$ and $200\times 200\times 16$, respectively. We further fixed the projection matrices under two resolution settings during testing to make the comparison. The model is trained using eight A10G GPUs with 24GB of memory, and it has been trained for 24 epochs on 3D semantic occupancy prediction tasks.
IV-B Dataset
The nuScenes dataset [14], a vast autonomous driving dataset, serves as the data source for our experiments. As the test set lacks semantic labels, we train our model on the training set and assess its performance using the validation set. We set the range for occupancy prediction as [-50, 50] meters for the X and Y axes and [-5, 3] meters for the Z axis. The input images have a $1600\times 900$ pixels resolution, while the final output occupancy is represented with a resolution of $256\times 256\times 32$ for the base version and $200\times 200\times 16$ for the small base version. We have conducted experiments on 3D semantic occupancy prediction tasks to provide quantitative results. The dense labels used for the 3D semantic occupancy prediction task are sourced from [39]. Additionally, we provide qualitative visualizations of the results for the 3D semantic occupancy prediction task.
IV-C Performance Evaluate Metrics
@@
IV-D Model Performance Analysis
We perform the task of multi-camera 3D semantic occupancy prediction on the nuScenes dataset. To evaluate the performance of our proposed InverseMatrixVT3D, we compare it with SOTA algorithms and present the results in Table I. All methods are trained using dense 3D occupancy ground truth labels obtained from [39]. Our model has two different settings for two resolutions: the InverseMatrixVT3D(Base) and InverseMatrixVT3D(Small Base), which includes the projection matrices generation process, and the InverseMatrixVT3D(Fix) and InverseMatrixVT3D(Small Fix), which fixes the projection matrices. Our model displays remarkable performance in both settings in predicting crucial dynamic objects on the road, including buses, cars, construction vehicles, and trucks. Additionally, our model excels in identifying diverse background objects, such as drivable surfaces, terrain, and vegetation. These results unequivocally highlight our model’s superior 3D world modelling capability compared to other SOTA algorithms.
Our model aggregates features for the final 3D volume based on sampling at specific sampling locations. Compared to the transformer-based approach, whose sampling locations can be adjusted based on the query vector, our approach’s sampling locations are fixed. Thus, when dealing with some small objects such as traffic cones and bicycles, our method aggregates fewer features from the image and is inferior to SurroundOcc in predicting those classes. However, our sampling density is significantly larger than the transformer-based approach, enabling our method to make good predictions on relatively large objects in the scene. In summary, our method sacrifices sampling flexibility to improve the sampling density. This underlying principle adjustment boosts model performance significantly in predicting relatively large objects in the scene.
Method Resolution 3DOcc mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [37] $200\times 200\times 16$ 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [60] $200\times 200\times 16$ 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [26] $200\times 200$ 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [38] $200\times 200\times 16$ 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 SurroundOcc [39] $200\times 200\times 16$ 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 InverseMatrixVT3D (Base) $256\times 256\times 32$ 26.03 26.15 9.76 42.27 41.89 17.83 14.74 22.88 10.30 21.78 36.90 50.82 21.15 24.45 26.76 16.86 32.03 InverseMatrixVT3D (Small Base) $200\times 200\times 16$ 25.95 26.13 15.19 42.65 43.92 11.42 20.82 22.68 13.45 22.33 36.04 47.47 16.26 21.47 25.33 16.37 33.73 InverseMatrixVT3D (Fix) $256\times 256\times 32$ 24.58 23.19 7.86 43.17 43.16 16.80 16.03 18.16 8.81 24.67 35.93 41.60 17.68 21.11 21.54 22.06 31.45 InverseMatrixVT3D (Small Fix) $200\times 200\times 16$ 22.68 20.91 12.06 38.52 35.53 10.53 21.49 20.12 12.36 20.49 30.31 36.23 14.60 19.25 22.78 15.89 31.76
TABLE I: 3D semantic occupancy prediction results on nuScenes validation set. All methods are trained with dense occupancy labels from [39].
IV-E Challenging Scenes Qualitative Analysis
Figure 6a presents a scene where multiple objects of different semantic classes overlap and crowd. Despite this complexity, our method accurately predicts each object’s semantic class and 3D spatial location. Furthermore, the edges between objects of different classes are delineated. Figure 6b exhibits the scene where a construction vehicle is severely obscured by a large truck while multiple traffic cones and barriers crowd and overlap. Even under such a complex scenario, our model can still infer the general occupancy status of the construction vehicle and generate a meticulous 3D occupancy prediction with distinct boundaries.
We further demonstrate our model’s powerful 3D modelling capability by presenting the prediction results in challenging rainy and night-time scenes, as shown in Figure 8. In each subfigure, the prediction result generated by our model is displayed on the bottom left, while the ground truth is shown on the bottom right. Regarding the rainy scenario depicted in Figure 7a, we observed an inconsistency in the ground truth labels. This discrepancy is attributed to the lidar’s poor performance during rainy weather due to the reflection effect. Consequently, the ground truth labels become inconsistent. However, our model’s prediction result is remarkably accurate. It not only fills in the missing information from the ground truth but also successfully detects a pedestrian crossing the road (indicated by the red circle), even though the person is located at a significant distance from the ego vehicle. Moreover, in the night-time scene illustrated in Figure 7b, our model effectively predicts the presence of a motorcycle (marked by the red circle) in the distance ahead despite the low ambient light conditions.
We also present scenes showcasing crowded multiple objects and rainy day crossroads in Figure 9. The left column displays the multi-camera images, the middle column shows the prediction results of our model, and the right column exhibits the ground truth labels. In the scenarios with crowded multiple objects represented in the first and second rows of Figure 9, our model accurately segments each object. Furthermore, in the rainy city intersection scenario depicted in the last row of Figure 9, our model successfully predicts the road structure even in the presence of missing information in the ground truth labels.
a Highly overlapping scene
b Severely obscured scene
@@ IV-E Challenging Scenes Qualitative Analysis
Figure 7: Challenging scenes qualitative analysis. Despite the presence of multiple objects that overlap or severely obscure each other, our method is capable of accurately predicting detailed occupancy and delineating their edges. Better viewed when zoomed in.
a Rainy scene
b Night-time scene
@@
Figure 8: Rainy and night-time scene. Despite challenging lighting conditions, our approach successfully predicts moving objects with high accuracy, even when they are far away from the ego vehicle. Better viewed when zoomed in.
a Multi-Camera Images
b Prediction Result
c Ground-Truth
Figure 9: Challenge scenes. Despite multiple different objects being crowded or missing information in the ground truth, our model can still accurately segment each object.
IV-F Model Efficiency
@@
Method Latency (s) ($\downarrow$) Memory (GB) ($\downarrow$) NeWCRFs [61] 1.07 14.5 MonoScene [37] 0.87 20.3 Adabins [62] 0.75 15.5 SurroundDepth [18] 0.73 12.4 SurroundOcc [39] 0.34 5.9 TPVFormer [38] 0.32 5.1 BEVFormer [26] 0.31 4.5 InverseMatrixVT3D (Base) 0.68 10.2 InverseMatrixVT3D (Small Base) 0.5 7.8 InverseMatrixVT3D (Fix) 0.45 4.88 InverseMatrixVT3D (Small Fix) 0.34 4.82
@@
TABLE II: Model efficiency comparison of different methods. The experiments are performed on a single RTX 3090 using six multi-camera images. For input image resolution, all methods adopt $1600\times 900$. $\downarrow$:the lower the better.
@@
Table II presents a comparison of the inference time and inference memory among different methods. The experiments are conducted on a single RTX 3090 using six multi-camera images. All methods adopt an image resolution of $1600\times 900$. The two base version models, which include the generation of projection matrices, have runtimes of 0.68 and 0.5 seconds for a single data sample. However, by fixing the projection matrices and removing the generation procedure, the fixed version model achieves exceptional real-time performance due to the utilization of projection matrices that facilitate 3D volume generation.
IV-G Ablation Study
IV-G1 Global Local Attention Fusion
@@
We perform an ablation study on the global-local attention fusion module, and the experiment results are presented in Table III. When we remove the fusion module and directly use the local 3D feature volume as the final 3D volume, we observe around 1.9% degradation in the model’s performance. This degradation could be attributed to the inability of the final 3D volume to capture long-range global semantic information.
Method 3DOcc mIoU w fusion module 23.19 w/o fusion module 21.32
@@
TABLE III: Ablation study for global-local attention fusion module
IV-G2 Multi-Scale Supervision
@@
We conducted an ablation study on the multi-scale supervision mechanism, and the experiment results are presented in Table IV. The upsample and skip connection structure remains in the model where the multi-scale supervision mechanism is removed, but we only utilize the highest-level 3D feature volume for supervision. The experimental results demonstrate that the coarse-to-fine supervision schema can enhance performance by approximately 2.2%.
Method 3DOcc mIoU w Multi-Scale Supervision 23.19 w/o Multi-Scale Supervision 21.07
TABLE IV: Ablation study for Multi-Scale Supervision
IV-G3 Dividing Schemas For Each Level
@@
We performed an ablation study on the division schemas for each level, and the experiment results are presented in Table V. We used different N settings for levels 1, 2, and 3 to generate sample points, and the experimental results aligned with our expectations. Specifically, when the divide setting N is larger, more sample points can be generated, leading to dense sampling and rich feature aggregation. We also observed a trend where a larger divide setting at higher levels has a more significant impact on the final model performance, indicating the increased importance of higher-level feature aggregation.
N setting of level 1,2,3 3DOcc mIoU $[3,4,5]$ 23.19 $[2,3,4]$ 20.01 $[1,2,3]$ 17.24 $[1,4,6]$ 19.33
@@
TABLE V: Ablation study of dividing schemas for each level
@@ V Conclusion
In this paper, we propose InverseMatrixVT3D, a method for vision-centric 3D semantic occupancy prediction. Our approach leverages predefined sample points for each scale of 3D volumes and constructs projection matrices to represent the fixed sampling process. Through matrix multiplication between multi-view feature maps and projection matrices in a multi-scale fashion, we generate local 3D feature volumes and global BEV features. These features are further merged using our proposed global-local fusion module, resulting in the final 3D volume at each level. Lastly, the 3D volumes at each level are upsampled and fused using a 3D deconvolution layer. Unlike other SOTA algorithms, our approach does not require depth estimation or transformer-based query, making the 3D volume generation process simple and efficient. Experimental results on nuScenes datasets demonstrate the superiority of our method in benchmark comparisons.

### v2

@@ IV-A Implementation Details
The InverseMatrixVT3D uses ResNet101-DCN [32] as a 2D backbone with a checkpoint from FCOS3D [33] to extract image features. The features of stage 1,2,3 of the backbone are fed to FPN [34], resulting in 3-level multi-scale image features. The network architecture comprises four levels ($L=4$), with no skip connection applied to the highest level. For the paths corresponding to levels 1, 2, and 3, we employ divided schemas of N=3, 4, and 5, respectively, to create sets of sample points. The AdamW optimizer with an initial learning rate of 5e-5 and weight decay of 0.01 is employed for optimization. The learning rate is decayed using a multi-step scheduler. Regarding data augmentation, random resize, rotation and flip operations are applied in the image space, following established practices for BEV-based 3D object detection [5, 6, 35, 7] and the compared methods [13, 14, 15, 25]. The predicted occupancy has a resolution of $200\times 200\times 16$ for full-scale evaluation. The model is trained to utilize eight A10 GPUs with 24GB of memory, and it has been trained for 2 days.
IV-B Loss Function
To train the model, we use focal loss [36], Lovasz-softmax loss [37] and scene-class affinity loss [25] to handle the significant sparsity of free space in the ground truth dataset. The final loss is composed of:
$Loss=L_{focal}+L_{lovasz}+L_{scal}^{geo}+L_{scal}^{sem}$ (5)
IV-C Dataset
The nuScenes dataset [1], a vast autonomous driving dataset, serves as the data source for our experiments. As the test set lacks semantic labels, we train our model on the training set and assess its performance using the validation set. We set the range for occupancy prediction as [-50, 50] meters for the X and Y axes and [-5, 3] meters for the Z axis. The input images have a $1600\times 900$ pixels resolution, while the final output occupancy is represented with a resolution of $200\times 200\times 16$ for the base version. We have conducted experiments on 3D semantic occupancy prediction tasks to provide quantitative results. The dense labels used for the 3D semantic occupancy prediction task are sourced from [14]. Additionally, we provide qualitative visualizations of the results for the 3D semantic occupancy prediction task.
To enhance the demonstration of the effectiveness of our approach, we conducted a monocular semantic scene completion experiment on the SemanticKITTI dataset [2] employing the left RGB camera. SemanticKITTI contains annotated outdoor LiDAR scans with 21 semantic labels. The input image resolution is $1241\times 376$, and the ground truth is voxelized into a grid of dimensions 256x256x32 with a voxel size of 0.2m. The evaluation of our model is performed on the validation set.
IV-D Performance Evaluate Metrics
@@
Refer to captionRefer to caption
@@
Figure 6: Challenging scenes qualitative analysis. Despite challenging lighting and weather conditions, our approach successfully predicts moving objects with high accuracy, even when they are far away from the ego vehicle. Better viewed when zoomed in.
IV-E Model Performance Analysis
We perform the task of multi-camera 3D semantic occupancy prediction on the nuScenes dataset and the monocular semantic scene completion task on the SemanticKITTI dataset. To evaluate the performance of our proposed InverseMatrixVT3D in each task, we compare it with other SOTA algorithms and present the results in Table I and II, respectively. In Table I, our model has three different settings: the InverseMatrixVT3D (Base), where the projection matrices generation process is included during training and evaluation, the InverseMatrixVT3D (Post-Fix), which involves fixing the projection matrices after the model finished training for evaluation purposes, and the InverseMatrixVT3D (Pre-Fix), which fixed projection matrices in advance and directly incorporating them into the training process.
In the context of the 3D semantic occupancy prediction task using the nuScenes dataset, our model demonstrates very competitive performance. It outperforms several transformer-query-based methods and ranks second on the benchmark. Compared to transformer-query-based methods, our model exhibits the best performance in detecting VRU on roadways, including bicycles, motorcycles, and pedestrians. Nevertheless, when dealing with background objects, like vegetation, manmade structures, and terrain, our model is inferior to transformer-query-based approaches. This may be attributed to the fact that many background objects are invisible, but due to the powerful generalization ability provided by the transformer, those transformer-query-based approaches can infer the background object’s semantic label and its locations to some extent. Another point worth mentioning is that although our model failed to beat SurroundOcc, the size of our model only has a total of 67M trainable parameters, which is substantially smaller than SurroundOcc’s 180M trainable parameters. Notably, the model’s performance significantly diminishes when utilizing fixed projection matrices after training due to noticeable projection errors. The projection error can be alleviated by employing fixed projection matrices in advance and directly incorporating them into the training process. In the realm of the monocular semantic scene completion task on the SemanticKITTI dataset, our model exhibits similar characteristics to those observed in the 3D semantic occupancy prediction task and has delivered very competitive performance in comparison with other SOTA algorithms. These results highlight our model’s superior 3D world modelling capability.
In general, our model aggregates features for the final 3D volume based on sampling at specific sampling locations. Compared to the transformer-based approach, whose sampling locations can be adjusted based on the query vector, our approach’s sampling locations are fixed. Nevertheless, our method boasts a significantly higher sampling density than the transformer-based approach, allowing for dense feature aggregation. In essence, our model prioritizes sampling density over sampling flexibility, which leads to competitive performance and a much smaller model size.
IV-F Challenging Scenes Qualitative Analysis
We demonstrate our model’s powerful 3D modelling capability and exceptional VRU detection performance by presenting the prediction results in challenging rainy and night-time scenes along with SurroundOcc prediction results and ground truth, as shown in Fig. 6. Regarding the rainy scenario depicted in Fig. 6 upper, we observed an inconsistency in the ground truth labels. This discrepancy is attributed to the lidar’s poor performance during rainy weather due to the reflection effect. Consequently, the ground truth labels become inconsistent. However, our model’s prediction result is remarkably accurate. It not only fills in the missing information from the ground truth but also successfully detects a pedestrian crossing the road (indicated by the red circle), even though the person is located at a significant distance from the ego vehicle. Moreover, in the night-time scene illustrated in Fig. 6 bottom, our model effectively predicts the presence of a motorcycle (marked by the red circle) in the distance ahead despite the low ambient light conditions.
IV-G Model Efficiency
@@
Method Latency (s) ($\downarrow$) Memory (GB) ($\downarrow$) NeWCRFs [38] 1.07 14.5 MonoScene [25] 0.87 20.3 Adabins [39] 0.75 15.5 SurroundDepth [3] 0.73 12.4 SurroundOcc [14] 0.34 5.9 TPVFormer [13] 0.32 5.1 BEVFormer [7] 0.31 4.5 InverseMatrixVT3D (Base) 0.5 5.2 InverseMatrixVT3D (Fix) 0.32 4.82
@@
TABLE III: Model efficiency comparison of different methods. The experiments are performed on a single RTX 3090 using six multi-camera images. For input image resolution, all methods adopt $1600\times 900$. $\downarrow$:the lower, the better.
@@
Table III compares the inference time and inference memory among different methods. The experiments are conducted on a single RTX 3090 using six multi-camera images. All methods adopt an image resolution of $1600\times 900$. Our base model, which includes the projection matrices generation, runs 0.5 seconds for a single data sample, but if we fixed the projection matrices, remove the projection matrices generation procedure. Due to the utilization of projection matrices that facilitate the 3D volume generation, our fixed version model achieves exceptional real-time performance.
IV-H Ablation Study
IV-H1 Global Local Attention Fusion
@@
We perform an ablation study on the global-local attention fusion module, and the experiment results are presented in Table IV. The experimental results confirm that the BEV feature and its associated refinement procedures are crucial in improving model performance. Without the BEV feature, the final 3D volume is unable to capture long-range global semantic information, ultimately leading to performance degradation.
[…]
@@
TABLE IV: Ablation study for global-local attention fusion module. Bott.ASPP: bottleneck aspp, Eff.Win.Atten: efficient window attention. $\uparrow$:the higher, the better.
IV-H2 Multi-Scale Mechanism
@@
We conducted an ablation study on the multi-scale mechanism, and the experiment results are presented in Table V. The experimental results demonstrate that the multi-level supervision and coarse-to-fine refinement structure play a vital role during training. Without these two structures, our model experiences a performance degradation of at least 5%.
[…]
TABLE V: Ablation study on multi-scale mechanism. Multi.Stru: multi-scale corse-to-fine refinement structure, Multi.Sup: multi-level supervision mechanism. $\uparrow$:the higher, the better.
IV-H3 Dividing Schemas For Each Level
@@
We performed an ablation study on the division schemas for each level, and the experiment results are presented in Table VI. We used different N settings for levels 1, 2, and 3 to generate sample points, and the experimental results aligned with our expectations. Specifically, when the divide setting N is larger, more sample points can be generated, leading to dense sampling and rich feature aggregation. We also observed a trend where a larger divide setting at higher levels significantly impacts the final model performance, indicating the increased importance of higher-level feature aggregation.
[…]
3 4 5 14.44%
2 3 4 8.45%
1 2 3 5.66%
1 4 6 7.73%
@@
TABLE VI: Ablation study of dividing schemas for each level. $\uparrow$:the higher, the better.
@@ V Conclusion
In this paper, we propose InverseMatrixVT3D, a vision-centric 3D semantic occupancy prediction method. Our approach leverages predefined sample points for each scale of 3D volumes and constructs projection matrices to represent the fixed sampling process. Through matrix multiplication between multi-view feature maps and projection matrices in a multi-scale fashion, we generate local 3D feature volumes and global BEV features. These features are merged using our proposed global-local fusion module, resulting in the final 3D volume at each level. Lastly, the 3D volumes at each level are upsampled and fused using a 3D deconvolution layer. Unlike other SOTA algorithms, our approach does not require depth estimation or transformer-based query, making the 3D volume generation process simple and efficient. Extensive experiments conducted on the nuScenes and SemanticKITTI datasets demonstrate that our method excels in its simplicity and effectiveness and achieves the best performance in detecting VRU for autonomous driving and road safety.